OpenStudy (anonymous):

Where do you break up the problem to have to solutions? Solve for x: 2|x - 3| + 1 = 7

OpenStudy (johnweldon1993):

After you get the absolute value by itself... Subtract 1 from both sides Divide both sides by 2 THEN you make it 2 equations

OpenStudy (anonymous):

How do you get it by itself, like the solution inside the bars?

OpenStudy (johnweldon1993):

Alright so we have \[\large 2|x - 3| + 1 = 7\] Subtract 1 from both sides \[\large 2|x - 3| = 6\] Divide both sides by 2 \[\large |x - 3| = 3\] Now you break it up into 2 equations... \[\large x - 3 = 3\] and \[\large x - 3 = -3\]

OpenStudy (anonymous):

Okay that confused me a bit? How do you know when to separate the equation?

OpenStudy (johnweldon1993):

So you see how I went through the steps? When I got to the step where the ONLY thing left on one side of the equation is the absolute value? |x - 3| = 3 ^ This is the only thing on this side of the equation... When that is the case...that is when you break it up

OpenStudy (anonymous):

oh okay, that makes sense now !!!! lol & would that be the answer or do i keep going?

OpenStudy (johnweldon1993):

Well That IS where you separate the equation... Do you need to solve for 'x' ?

OpenStudy (anonymous):

yes i need to solve for X. I got 0 on one of the solutions

OpenStudy (johnweldon1993):

0 is one of the solutions \[\large x - 3 = -3\] But what about \[\large x - 3= 3\]

OpenStudy (anonymous):

it would be 0 again:?

OpenStudy (johnweldon1993):

\[\large 0 - 3 \ne 3\]

OpenStudy (anonymous):

6?

OpenStudy (johnweldon1993):

There we go! So your answers are both 0 and 6

OpenStudy (anonymous):

Thank You!!

OpenStudy (johnweldon1993):

No problem!